OpenStudy (anonymous):

Find the radius of a sphere with a surface area of 804cm2 @iGreen

OpenStudy (igreen):

\(SA_{sphere} = 4 \pi r^2\) Plug in what we know: \(804 = 4(3.14) r^2\) Multiply 4 * 3.14

OpenStudy (anonymous):

12.56 @iGreen

OpenStudy (igreen):

Yes, so we have: 804 = 12.56r^2 Divide 12.56 to both sides, what's 804 / 12.56?

OpenStudy (anonymous):

64 (rounded) @iGreen

OpenStudy (igreen):

Yes, so we have: 64 = r^2 Find the square root of both sides, what's the square root of 64?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks again! @iGreen

OpenStudy (igreen):

Yep, that's our answer.

OpenStudy (anonymous):

thank you!